OpenStudy (anonymous):

How do I express: 1/2(r+2)^2 = 1/3(2r-1)^2 in simplest radical form?

OpenStudy (anonymous):

\[\left(\begin{matrix}1 \\ 2\end{matrix}\right)(r + 2)^{2} = \left(\begin{matrix}1 \\ 3\end{matrix}\right)(2r-1)^{2}\]

OpenStudy (anonymous):

you don't.

OpenStudy (anonymous):

yes, I do. The answer is -7+3\[\sqrt{5}/2\]

OpenStudy (anonymous):

congrats